OpenStudy (anonymous):

What is the ratio 4 to 60 expressed as a fraction in simplest form?

OpenStudy (anonymous):

\[\frac{4}{60}\]

OpenStudy (anonymous):

Do you have any answer or have you tried it yet?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

Ratio of \(m\) to \(n\) is given as: \[\frac{m}{n}\]

OpenStudy (anonymous):

Now tell me, out of 4 and 60, can you think of any number which can divide both 4 and 60 at the same time?

OpenStudy (anonymous):

What's 60 / 4? @CuteGirl903k

OpenStudy (anonymous):

15

OpenStudy (anonymous):

Yep, and 4 / 4 = 1, so that gives us \(\dfrac{1}{15}\).

OpenStudy (anonymous):

I can show you more clearly. :)

OpenStudy (anonymous):

\[\frac{\cancel{4}^2}{\cancel{60}^{30}} = \frac{2}{30} = \frac{\cancel{2}^1}{\cancel{30}^{15}} = \color{blue}{\frac{1}{15}}\]

OpenStudy (anonymous):

Ok!!! Thanks everyone!!!!:D